OpenStudy (anonymous):

Write each equation in standard form. y + 6 = (x + 4) Select one: a. x – y = 10 b. y = x – 2 c. x + y = –2 d. x – y = 2

OpenStudy (anonymous):

-x+y=-2

OpenStudy (anonymous):

x-y-2=0 is standard form of linear eq. ax+by+c=0 this can be sometimes written as ax+by = c, so desired standard eq is x-y =2